OpenStudy (oayala):

( 2 x 10^13 )^-2 in scientific notation?

OpenStudy (seratul):

2*10^13 is equal to 20000000000000 20000000000000^-2 is 200000000000 200000000000 in scientific notation is 2*10^11

OpenStudy (robtobey2):

\[\frac{1}{\left(2\ 10^{13}\right)^2}=\frac{1}{4\ 10^{26}}=\frac{1}{400000000000000000000000000} \]